OpenStudy (anonymous):

add or subtract -4x/x+7 - 8/x+7

OpenStudy (anonymous):

i don't get the question. you can say either way, using, people would say subtracting, but we can also say that we're adding the opposite. for the numerator (-4x -8) is the same as (-4 + -8) it's your pick really. Lower level math, i think they tend to go with -4x - 8. But sometimes, the other way can help you organize better. Good luck, and let me know if i did any error :)

OpenStudy (anonymous):

yeah it's (-4x-8) but how did you get that answer?

OpenStudy (anonymous):

i mean it's -4x-8/x+7 but i wanna know how i can say i got that answer because my teacher want me to explain how i got it, i just guesses she said it was right but i need to explain

OpenStudy (anonymous):

common denominator, just add up the numerator :) imagine your using 4+5, it can also be written as 4/1 +5/1 right? we add it up, and it becomes 9/1 = 9. However, when we have 1/2 +1/2, we add them up like this: (1+1)/2 = 2/2 = 1. again, it makes sense that if you have 1/2 a pie, then another 1/2 of the same pie. When you put 2 together, you have a full pie! Hope you understand the concept. Good luck. Note: don't try to add this way when the denominator (bottom part) are different though, otherwise it will be a disaster!

OpenStudy (anonymous):

k thanks